How to find root of 8+3i in complex numbers???????????????????????

Grade:11

1 Answers

Amogh Dasture
8 years ago
Let x+iy=root of 8+3i
 
(x+iy)2=8+3i
 
x2-y2+2xyi=8+3i 
 
On comparing, x2-y2=8 and 2xy=3 ie x=3/(2y)
 
9/(4y2) – y2=8
 
Solve and get the values of x and y 
 
z=x+iy